OpenStudy (anonymous):

At a party, everyone shook hands with everybody else. There were 66 handshakes. How many people were at the party?

OpenStudy (anonymous):

math help!

OpenStudy (anonymous):

@thisgirl9925 please message me!

OpenStudy (anonymous):

Wouldn't you divide by 2? If there are 66 people, and two people have to shake hands... maybe 33, but not sure.

OpenStudy (anonymous):

thanks for the help but im not sure thats it

OpenStudy (anonymous):

@thisgirl9925 :/

OpenStudy (anonymous):

It wouldn't be divided by 2 for sure

OpenStudy (anonymous):

this questions hurts my head

OpenStudy (anonymous):

Lets say that there's 4 people at a party, and we name them A, B, C, and D

OpenStudy (anonymous):

All the possible hand shakes are AB, AC, AD, BC, BD, and CD

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

But we have to do that with 66 somehow

OpenStudy (anonymous):

@SarahEZZMcK

OpenStudy (anonymous):

Ok I got something

OpenStudy (anonymous):

It looks like this is something called a triangular number sequence: http://www.mathsisfun.com/algebra/triangular-numbers.html

OpenStudy (anonymous):

And we should be able to use this formula to solve: \[x _{n} = n(n+1)/2\]

OpenStudy (anonymous):

ummm idek how to do that

OpenStudy (nathan917):

12 people. n=11 (when there are 12 people, there will be only 11 hand shake) For checking it=n(n+1)/2=11*12/2=6*11=66 shakehands

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

@thisgirl9925 unblock me pleasse!